OpenStudy (anonymous):

solve 3(c+2b)=d for c

OpenStudy (anonymous):

is that the complete question?

OpenStudy (anonymous):

thats what i got is for a review on literal equations

OpenStudy (anonymous):

solve 3(c+2b)=d for c divide 3 on both sides, c+2b=d/3 minus 2b on both sides, c=(d/3)-2b

OpenStudy (anonymous):

thank you!